a die and a coin are tossed. what is the probability fo tossing a 5 and a head?
Answer by stanbon(75887)

a die and a coin are tossed. what is the probability fo tossing a 5 and a head?
-----------------
These are independent events, so you multiply to get the joint probability.
P(5 and head) = P(5)*P(head) = 1/6*1/2 = 1/12
======================
Cheers,
Stan H.